OpenStudy (anonymous):

When asked to balance the equation: C2H6 + O2 = CO2 + H2O the following suggestions were made: C2H6 + 5O2 = 2CO2 + 3H2O C2H6 + 5O = 2CO + 3H2O 2C2H6 + 7O2 = 4CO2 + 6H2O Which answer is correct and what is wrong with the others?

OpenStudy (mertsj):

In the first answer, how many carbon atoms on the left side?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

how many carbon atoms on the right side?

OpenStudy (photon336):

the number of each atom must be equal on each side of the equation remember law of conservation of mass

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so..which one of these answers is correct?

OpenStudy (photon336):

@1stwilight before I help you out can you tell me which one you think it is and why?

OpenStudy (photon336):

@15Twilight

OpenStudy (anonymous):

It looks like all of them are correct

OpenStudy (anonymous):

but I know that's not right

OpenStudy (photon336):

@15Twilight follow this to see whether an equation is balanced make a chart. and count up the number of each atom in both the reactants and the products. C2H6 + O2 = CO2 + H2O  based of of this is the equation balanced?

OpenStudy (anonymous):

no

OpenStudy (photon336):

good. for your reference always do that in the future

OpenStudy (photon336):

we need to balance this now, this takes time and there's no real i guess strategy you just have to try your best to put coefficients infront of the reactants+ products to balance them.

OpenStudy (photon336):

C2H6 + 4O2 = 2CO2 + 3H2O

OpenStudy (anonymous):

2 atoms of carbon 6 atoms of hydrogen, 8 atoms of oxygen products: 2 atoms of carbon 4 atoms of oxygen, and 6 atoms of hydrogen ..am I right?